John currently has enough money to buy 45 books. If the cost of each book was 10 cents less, John could buy 5 more books. How much money does John have to spend on books?

Answer by richwmiller(17219)

45x=y
50(x-10)=y
45x=50(x-10)
45x=50x-500
500=5x
100=x
45*100=y
$45.00 cents =y
y=$45.00 to spend on books
check
45*100=4500
50*90=4500
ok